# LLMs are not Zero-Shot Reasoners for Biomedical Information Extraction

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly adopted for applications in healthcare, reaching the performance of domain experts on tasks such as question answering and document summarisation. Despite their success on these tasks, it is unclear how well LLMs perform on tasks that are traditionally pursued in the biomedical domain, such as structured information extration. To breach this gap, in this paper, we systematically benchmark LLM performance in Medical Classification and Named Entity Recognition (NER) tasks. We aim to disentangle the contribution of different factors to the performance, particularly the impact of LLMs’ task knowledge and reasoning capabilities, their (parametric) domain knowledge, and addition of external knowledge. To this end we evaluate various open LLMs—including BioMistral and Llama-2 models—on a diverse set of biomedical datasets, using standard prompting, Chain-of-Thought (CoT) and Self-Consistency based reasoning as well as Retrieval-Augmented Generation (RAG) with PubMed and Wikipedia corpora. Counter-intuitively, our results reveal that standard prompting consistently outperforms more complex techniques across both tasks, laying bare the limitations in the current application of CoT, self-consistency and RAG in the biomedical domain. Our findings suggest that advanced prompting methods developed for knowledge- or reasoning-intensive tasks, such as CoT or RAG, are not easily portable to biomedical tasks where precise structured outputs are required. This highlights the need for more effective integration of external knowledge and reasoning mechanisms in LLMs to enhance their performance in real-world biomedical applications.
Introduction
The success of Large Language Models (LLMs) promises to reshape the landscape of AI healthcare applications, especially for scenarios relying on Question Answering (Budler, Gosak, and Stiglic 2023; Subramanian et al. 2024), summarisation (Schlegel et al. 2023) and extracting insights from unstructured patient-generated health data (Li et al. 2023). While considerable progress has been made in leveraging LLMs for tasks requiring free-text outputs, much of the focus has been on optimizing the parametric knowledge—the information stored in the model’s weights and learned during training. Recent works explore methods such as fine-tuning on task-specific data and in-context learning (ICL) and reporting significant improvements in model performance.
@@ Introduction
However, these approaches primarily enhance the models’ internal knowledge representation. As such, they rely on readily available data for the structured tasks at hand, be it in form of training sets for task-specific fine-tuning (Abburi et al. 2023), or for selecting good-quality representative few-shot examples for ICL (Zhang et al. 2024; Gutierrez et al. 2022). In the biomedical domain, such resources for structured prediction tasks are typically not available, as requirements might arise ad-hoc—for example when researchers need to process a set of medical records to find patients satisfying inclusion criteria for a clinical trial (Jullien et al. 2023) (e.g., whether they’re a smoker). But even for well established tasks, such as medication name extraction, for which resources exist (Wei et al. 2020), these resources often prove to be insufficient in a practical context, due to the domain shift between public resources and internal hospital data (Hadi et al. 2023). Therefore, solely training-set reliant improvements parametric knowledge of LLMs as driver of performance for structured prediction tasks is often infeasible and approaches need to be able to perform well in zero-shot scenarios. Despite this, the literature currently lacks a systematic investigation of other crucial aspects of knowledge utilization.
@@
In order to address this research gap, we first postulate that the performance of LLMs in medical reasoning and information extraction tasks in “true” zero-shot setting111by “true zero-shot” we refer to the scenario where no examples are available to solve the task and no information beyond the labels and their semantically meaningful names is made available to the model (Lampert, Nickisch, and Harmeling 2014). hinges on three distinct categories of knowledge:
  • Parametric Knowledge: The inherent knowledge embedded within the model’s parameters.
  • Task Knowledge: The model’s ability to reason about the specific task, including understanding relevant labels and the context of the task.
  • External Knowledge: Additional information and context retrieved to supplement the model’s understanding and decision-making process.
@@
Research in evaluating these aspects specifically in the medical domain (Nori et al. 2023; Subramanian et al. 2024) is being conducted vividly, but these works have mostly focused on knowledge-intensive prerequisite tasks, such as Multiple-Choice Question Answering. While useful to evaluate the medical knowledge of LLMs, they do not address the question of the medical capabilities of LLMs to succeed on tasks that are more reflective of real applications, such as medical text classification or information extraction. As such, it is necessary to evaluate, whether advancements derived from methods that enhance performance, such as (zero-shot) Chain-of-Thought (CoT) reasoning (Wei et al. 2022; Wang and Zhou 2024), self-consistency (Wang et al. 2022) and Retrieval-augmented Generation (RAG) (Li et al. 2024) carry over to such structured prediction tasks.
Moreover, these studies often employ large, commercial models like ChatGPT (Biswas 2023) or GPT-4 (OpenAI 2023) which present significant challenges in real-world applications due to their computational cost and privacy concerns associated with sending sensitive data to third-party APIs. Furthermore, there is a growing concern regarding the reliability of LLMs in medical applications, as even the most powerful models are prone to generating hallucinations, compromising the truthfulness of the outputs. Although constrained generations have shown promise in mitigating these issues, their application in medical information extraction tasks has been limited.
Thus, there are three problems that currently inhibit our understanding of the capabilities of LLMs on structured prediction tasks in the medical domain, and, as a consequence, their improvement: (i) Existing approaches to structured prediction tasks in the medical domain typically enhance parametric knowledge and rely on the availability of training sets, which might not be realistic; (ii) “True zero-shot” studies and methods to improve performance in such settings are mostly carried out surrogate tasks such as Question Answering and whether they can be adapted to structured prediction tasks on is unknown; (iii) Advancements are typically reported on large-scale, proprietary LLMs which might be unusable due to privacy concerns and inaccessibility to logits for constrained decoding.
In this paper, we aim to address these gaps by systematically benchmarking the performance of LLMs in medical classification and NER tasks as a representative selection of structured prediction tasks. We focus on assessing the impact of task knowledge and external knowledge while maintaining the parametric knowledge at a reasonable yet static level. Our approach involves exploring a range of techniques, including CoT reasoning, Retrieval-Augmented Generation (RAG), and constrained generation, which have not been extensively applied in these settings. By providing a comprehensive evaluation of these methods, we seek to offer new insights into the practical deployment of LLMs in the medical domain, highlighting both the challenges and potential solutions.
@@
To summarise, this paper makes the following novel contributions: First, to our knowledge, we present the first comprehensive benchmark for LLMs in medical classification and Named Entity Recognition (NER) tasks, providing a systematic evaluation of their information extraction performance in these critical structured prediction tasks within the medical domain. Second, we investigate the impact of various knowledge enhancement techniques, including Chain of Thought (CoT) reasoning, Self-Consistency, Retrieval-Augmented Generation (RAG), and constrained generation, which have not been extensively explored in medical information extraction settings. Notably, we demonstrate that parametric knowledge capacity, i.e., model size, is a primary and often sole driver of performance in zero-shot settings, offering insights into the limitations and potential of current LLM architectures.
Related Work
@@
We briefly survey the existing benchmarking literature in the medical domain, outlining the lack of studies focusing on structured prediction tasks. Furthermore, we cover recent prompting techniques that were proposed to elicit reasoning in LLMs, and augment their domain knowledge, either by better tapping into their parametric knowledge or by explicitly providing them with relevant external context. Notably, we omit approaches that rely on existence of training sets, such as few-shot prompting (Wang et al. 2023) or model fine-tuning, as one of the key challenges in the medical domain is the lack of annotated task data, due to privacy concerns over sharing medical records. Instead, as outlines in the introduction, we focus on “true” zero-shot capabilities of LLMs.
Existing LLMs Benchmarks: With the rising popularity of LLMs, many works evaluated their performance in the biomedical and clinical domains. These works typically focus on evaluating domain-knowledge by means of Question Answering (Singhal et al. 2023; Harris 2023; Subramanian et al. 2024), or focus directly on possible application scenarios, such as summarisation (Li et al. 2023; Yim et al. 2023) or clinical coding (Kaur, Ginige, and Obst 2023). Many works combine these two directions in an effort to provide more comprehensive benchmarks (Srivastava et al. 2024; Xiong et al. 2024; Feng et al. 2024; Chen et al. 2020; Manes et al. 2024). However, many of these works overlook the wealth of existing literature and plethora of available resources for traditional structured prediction tasks in the biomedical domain, such as document classification, entity recognition and linking and event and relation extraction (e.g., Pyysalo et al. (2007; 2012) to name a few). Fries et al. (2022) have provided a comprehensive and unified collection of these resources, however their work prioritises reportage of the resource collection over benchmarking results. Their preliminary evaluations suggest that their evaluated pre-LLM era models barely surpass the random guess baseline in the zero-shot setting. We build upon their work by providing a detailed analysis to what extent approaches to enhance reasoning and knowledge in LLMs help to challenge this status quo.
@@
Reasoning- and Knowledge-enhancing approaches: Current work attempts to improve the performance of LLMs from different knowledge utilization perspectives. One of the obvious methods is full parameter domain-specific pre-training (Xie et al. 2024). For example, Chen et al. (2023) propose the largest medical foundation model, trained on both biomedical and clinical data, up to 70B. Bolton et al. (2024), on the other hand, believe larger LLMs are computationally expensive to run, proposing a 2.7B LLM specific for biomedical NLP tasks. When fine-tuned, the relatively small model compete with larger LLMs. In our study, we compare domain-generalist models with those adapted to the medical domain. Since full parameter tuning is costly, many works focus on domain knowledge adaptation by pre-training (Shi et al. 2024; Song et al. 2024) or instruction tuning (Willard and Louf 2023) with adapters. Training-free approaches encompass chain-of-thought (CoT) (Wei et al. 2022; Jeong et al. 2024), self-consistency(Wang et al. 2022), Concerned with lack of grounding resulting in hallucination, recent work introduce RAG methods (Li et al. 2024; Wang et al. 2024b; Yu et al. 2023; Munnangi et al. 2024; Wang et al. 2024a; Soong et al. 2023). However, most of these efforts have focused on performance in a particular knowledge paradigm and have lacked a systematic assessment of how performance on structured prediction, which we address in our study.
Methodology
Our methodology is designed to answer the following two research questions: “How well do LLMs perform on structured prediction tasks?” and “To what extent can approaches that enhance task and external knowledge improve their performance?” To answer the first research question, we benchmark a representative sample of LLMs on a large collection of biomedical text classification and NER datasets. More specifically, we choose the task of Medical Text Classification and NER as representative structured predictions tasks. We focus on the “true” zero shot setting, since, as discussed before, this allows us to establish the level of models’ original parametric knowledge, which is desirable as it more closely reflects real-world application scenarios, because annotated training data for such tasks in the biomedical domain is usually not available due to the ad-hoc nature of task requirements and privacy constraints of medical records. Thus improving parametric knowledge is often infeasible in practice. To answer the second question, we compare their zero-shot performance to various methods that aim to enhance task knowledge and external knowledge, while keeping the parametric knowledge static.
Datasets
Since we evaluate different prompting techniques, we restrict the choice of tasks to those where the number of possible labels is small enough to fit in the evaluated LLMs’ context window. We restrict the number of labels to ten and the mean length of the input documents to at most 2048 tokens. This leaves us with 14 different classification datasets from the BigBio collection222for the GAD dataset, we only select 1 fold out of the 10 available, as the folds feature the same task for different data, unlike other datasets. We also skipped the Chinese subset of meddialog as we had difficulties loading the dataset. For the NER task, we sample 12 datasets from the pool of those that satisfy the criteria. The resulting dataset sample features four non-English datasets and six non-public classification datasets, which allows us to investigate whether LLMs perform better on minority languages or on data that is less likely to be found in public pre-training corpora. We run the evaluation on the official test-set split where available, otherwise we consider the full dataset. For datasets with more than 500 instances, we sample 500 random but fixed instances to speed up the experiments. Overall, our selection spans English and non-english source data, publicly available and private datasets, and various domains such as scientific papers, medical notes and social media. The overview of the datasets follows below, with full details to be found in the technical appendix.
Classification.
The datasets used for classification tasks include both single-label and multi-label datasets, covering a wide range of biomedical and clinical domains. For single-label classification, the GAD dataset focuses on identifying associations between genes and diseases (Bravo et al. 2015), while the GEO dataset is concerned with classifying microarray, transcriptomics, and single-cell experiments from the Gene Expression Omnibus (GEO) database (Elucidata 2022). The MedDialog dataset aims to classify dialogue snippets as either being said by a doctor or a patient (Chen et al. 2020). Furthermore, the CZIDrsm dataset has several subsets, including one for classifying research articles based on aspects of disease research (CZIBase), and others for identifying whether a paper describes substantive research into Quality of Life (CZIQoL) or is a natural history study (CZINatHist).
In multi-label classification, the LitCovid dataset is used for the classification of COVID-19-related articles (Chen et al. 2021). The CAS and ESSAI datasets are utilized for identify negation and uncertainty clinical cases from French-speaking countries (Grabar, Claveau, and Dalloux 2018). The NTCIR13 datasets include subsets for disease classification of tweets in Japanese (*-Ja), English (*-En), and Chinese (*-Zh) (Iso et al. 2017). Additionally, the PsyTAR dataset is used for sentence classification of various drug-related effects, such as Adverse Drug Reactions (ADR) and Withdrawal Symptoms (WDs) (Zolnoori et al. 2019), while the SciCite dataset is used for citation intent classification based on the context within computer science and biomedical domains (Cohan et al. 2019).
NER.
The datasets for Named Entity Recognition (NER) tasks are similarly divided into entity recognition (single entity type) and classification (multiple entity types). In the single-type category, the GeneTag dataset is used for gene/protein NER, with two annotation versions: the original GeneTag-G and the corrected GeneTag-C (Tanabe et al. 2005). Additionally, the GENIA-PPI dataset focuses on protein-protein interactions or gene regulatory relations within the GENIA corpus, capturing primarily static relations (Pyysalo et al. 2009; Hoehndorf et al. 2010; Ohta et al. 2010).
The multiple-type NER datasets encompass various complex biomedical tasks. The AnEm dataset targets anatomical entity recognition (Ohta et al. 2012), while the BioInfer dataset focuses on recognizing proteins, genes, and RNA entities (Pyysalo et al. 2007). The Genia-EE dataset is used for the GENIA Event corpus (Kim et al. 2009), and the BioNLP11-REL dataset is employed for extracting part-of relations between genes/proteins and associated entities (Pyysalo, Ohta, and Tsujii 2011). Furthermore, the BioNLP-13-CG dataset is used for Cancer Genetics (CG) information extraction, focusing on recognizing events represented as structured n-ary associations of given physical entities (Pyysalo, Ohta, and Ananiadou 2013). The BioNLP-13-GRO dataset aims to populate the Gene Regulation Ontology with events and relations (Kim et al. 2013), and the BioNLP-13-PC dataset is used for the automatic extraction of biomolecular reactions from text (Ohta et al. 2013). Lastly, the PICO dataset deals with recognizing (P)articipants, (I)nterventions, and (O)utcomes (Nye et al. 2018), and the MLEE dataset is used for event extraction related to angiogenesis (Pyysalo et al. 2012).
Models
For our experiments, we employed two instruction-tuned variants of the Llama-2 model—7B and 70B—both (Touvron et al. 2023), alongside the BioMistral-7B model (Labrak et al. 2024) which was further pre-trained on the biomedical domain. Since we make use of constrained generation to generate model outputs and guide the models decoding process, we retrict the evaluation to open source models since this process is not possible for proprietary models such as GPT-4.
@@
Standard prompting was used as a baseline for both the Classification as well as the NER tasks. Chain-of-thought reasoning (Wei et al. 2022) has been shown to improve performance, particularly in QA and logical reasoning tasks. Thus, we also ran experiments with chain-of-thought reasoning to measure its impact on model performance. For the NER task, we adapted a more guided, two-stage approach (Shen et al. 2021) to implement a novel chain-of-thought reasoning approach. Here, The first stage involves inducing a generic entity name from a datasets’ known entity labels—e.g., “Bodypart” for the NER labels describing different bodyparts—and then labelling the input document with that generic entity type. In the second stage all entities labelled in this way are further disambiguated with their respective fine-grained dataset NER labels. Retrieval Augmented Generation (Lewis et al. 2020) has been established as an effective technique to improve model performance by introducing relevant non-parameteric knowledge to models and thus grounding the generated outputs to factual information. Xiong et al. (2024) conducted a systematic study of RAG on medical QA, and we incorporate their findings into our study. We used PubMed abstracts (Sanyal, Bhowmick, and Das 2021) and Wikipedia articles as knowledge corpora, because Xiong et al.’s (2024) experiments found that using PubMed improved performance over non RAG techniques, while using Wikipedia reduced performance in medical QA tasks. Our goal was to evaluate whether the same holds true for structured prediction tasks as well. For the RAG module, we made use of FAISS (Douze et al. 2024; Johnson, Douze, and Jégou 2019), which allows retrieval of most similar documents based on semantic similarity, where we used the all-MiniLM-L6-v2 sentence transformers (Reimers and Gurevych 2019) model for embedding input documents and corpora. For each experiment, the number of retrieved documents was computed based on the maximum possible documents which could be used without exceeding the token limit of the model.
Self-consistency, proposed by Wang et al. (2022), improves chain-of-thought reasoning of LLMs by sampling reasoning paths for a given problem, followed by a majority vote for the final answer. We also conduct a set of experiments employing self-consistency to investigate whether such improvements can be observed on structured prediction tasks in the medical domain as well. For classification tasks, self consistency was employed to generate multiple reasoning chains for the given problem, followed by answer extraction from each reasoning chain and majority voting to select the final answer. For NER tasks, since we follow the two-stage approach, self-consistency was employed in both stages. Multiple general entity labels were generated in the first stage, and entities were extracted for each such label. In the second stage, self consistency was again used for the entity selection phase as well as the entity label determination step. Majority voting was utilised in final label or class selection in each case (Xie et al. 2023).
Constrained decoding in LLMs (Willard and Louf 2023) was used to ensure structured information extraction and text generation. This allowed us to evaluate the LLMs for the task at hand without the added variability due to the aleatoric uncertainties brought about by the probabilistic language generation fundamental to the architectures of the models. More specifically, for classification tasks, we ensured the presense of at least one label in the generated outputs. For NER we restricted the generation of spans occurring in text in the first step, and in the second step, for each of the spans we restricted the generation to any of the possible labels. This is also one of the reasons why we opted against evaluating API-based closed-source LLMs [^3] The other reason being their intransparancy with regard to training data, which violates our “true” zero-shot setting., as in our initial experiments the hallucinations in generated outputs created problems with reliably parsing the structured outputs.
Evaluation Results
Overview of results

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly adopted for applications in healthcare, reaching the performance of domain experts on tasks such as question answering and document summarisation. Despite their success on these tasks, it is unclear how well LLMs perform on tasks that are traditionally pursued in the biomedical domain, such as structured information extraction. To bridge this gap, in this paper, we systematically benchmark LLM performance in Medical Classification and Named Entity Recognition (NER) tasks. We aim to disentangle the contribution of different factors to the performance, particularly the impact of LLMs’ task knowledge and reasoning capabilities, their (parametric) domain knowledge, and addition of external knowledge. To this end, we evaluate various open LLMs—including BioMistral and Llama-2 models—on a diverse set of biomedical datasets, using standard prompting, Chain-of-Thought (CoT) and Self-Consistency based reasoning as well as Retrieval-Augmented Generation (RAG) with PubMed and Wikipedia corpora. Counter-intuitively, our results reveal that standard prompting consistently outperforms more complex techniques across both tasks, laying bare the limitations in the current application of CoT, self-consistency and RAG in the biomedical domain. Our findings suggest that advanced prompting methods developed for knowledge- or reasoning-intensive tasks, such as CoT or RAG, are not easily portable to biomedical tasks where precise structured outputs are required. This highlights the need for more effective integration of external knowledge and reasoning mechanisms in LLMs to enhance their performance in real-world biomedical applications.
LLMs are not Zero-Shot Reasoners for Biomedical Information Extraction
Aishik Nagar1  Viktor Schlegel2,3  Thanh-Tung Nguyen  Hao Li3 Yuping Wu3Kuluhan Binici4Stefan Winkler4 1ASUS Intelligent Cloud Services (AICS), Singapore 2Imperial College London, Imperial Global Singapore 3University of Manchester, United Kingdom 4National University of Singapore, Singapore Correspondence to: aishik_nagar@asus.com
1 Introduction
@@
The success of Large Language Models (LLMs) is reshaping AI healthcare applications, particularly in Question Answering Budler et al. (2023); Subramanian et al. (2024), summarization Van Veen et al. (2024); Schlegel et al. (2023); Nagar et al. (2024), and extracting insights from unstructured patient-generated health data Li et al. (2023). While advancements in fine-tuning and in-context learning (ICL) have improved LLM performance, these rely on readily available structured training data Abburi et al. (2023); Zhang et al. (2024); Gutierrez et al. (2022). However, in biomedical contexts, such resources are often unavailable due to domain shifts Hadi et al. (2023) or ad-hoc requirements —for example when researchers need to process a set of medical records to find patients satisfying inclusion criteria for a clinical trial  Jullien et al. (2023); Hadi et al. (2023) (e.g., whether they’re a smoker). This limits the effectiveness of parametric knowledge improvements in LLMs, necessitating strong zero-shot capabilities for structured prediction tasks such as biomedical classification and Named Entity Recognition. Despite this, the literature currently lacks a systematic investigation of other crucial aspects of knowledge utilization in zero-shot performance of LLMs in such tasks.
@@
In order to address this research gap, we first postulate that LLM performance in true zero-shot settings—where only task labels and their meaningful names are provided Lampert et al. (2014)—hinges on three categories of knowledge: (a) Parametric Knowledge: Information embedded in model weights; (b) Task Knowledge: Understanding of task-specific labels and context; (c) External Knowledge: Additional retrieved context to supplement decision-making.
@@
Existing research evaluating these factors in LLMs for the medical domain focus on knowledge-intensive tasks like Multiple-Choice QA Nori et al. (2023); Subramanian et al. (2024), but their capabilities in structured prediction tasks, such as medical classification and information extraction, remain underexplored. Additionally, techniques like zero-shot Chain-of-Thought (CoT) reasoning Wei et al. (2022); Wang and Zhou (2024), self-consistency Wang et al. (2022), and Retrieval-Augmented Generation (RAG) Li et al. (2024) require systematic evaluation in these contexts.
Additionally, evaluations often focus on proprietary models like ChatGPT Biswas (2023) or GPT-4 OpenAI (2023), which pose challenges due to computational cost, privacy concerns, and inaccessibility for techniques like constrained decoding. Despite the growing concerns regarding reliability of LLMs in medical applications, techniques like constrained decoding which have shown promise in mitigating LLM hallucinations have not been systematically applied to medical information extraction or classification.
Thus, four key issues currently hinder progress: (i) Reliance on training sets and parametric knowledge for structured prediction, which may be unrealistic; (ii) Lack of true zero-shot evaluations for structured tasks beyond surrogate QA; (iii) Dependence on large-scale, proprietary LLMs, limiting practical deployment. (iv) Lack of a systematic analysis of impact of latest techniques such as Chain-of-thought reasoning, RAG and constrained generation in medical structured prediction tasks.
This paper systematically benchmarks LLMs in medical classification and Named Entity Recognition (NER), assessing task and external knowledge while controlling parametric knowledge. We evaluate CoT reasoning, RAG, and constrained generation, offering insights into their applicability.
@@
First, we present the first comprehensive benchmark of task and external knowledge adaptation for LLMs in medical structured prediction tasks. Second, we analyze the impact of knowledge enhancement techniques, including CoT, self-consistency, RAG, and constrained generation. Third, we demonstrate that parametric knowledge capacity, i.e., model size, is the primary driver of zero-shot performance, highlighting the limitations and potential of current LLM architectures.
2 Related Work
@@
We survey the existing benchmarking literature for the medical domain in the appendix section A, outlining the lack of studies focusing on structured prediction tasks. Furthermore, we cover recent prompting techniques that were proposed to elicit reasoning in LLMs, and augment their domain knowledge, either by better tapping into their parametric knowledge or by explicitly providing them with relevant external context. Notably, we omit approaches that rely on existence of training sets, such as few-shot prompting Wang et al. (2023) or model fine-tuning, as one of the key challenges in the medical domain is the lack of annotated task data, due to privacy concerns over sharing medical records. Instead, as outlines in the introduction, we focus on “true” zero-shot capabilities of LLMs.
@@
Reasoning- and Knowledge-enhancing approaches: Current work attempts to improve the performance of LLMs from different knowledge utilization perspectives. One of the obvious methods is full parameter domain-specific pre-training Xie et al. (2024). For example, Chen et al. (2023) propose the largest medical foundation model, trained on both biomedical and clinical data, up to 70B. Bolton et al. (2024), on the other hand, believe larger LLMs are computationally expensive to run, proposing a 2.7B LLM specific for biomedical NLP tasks. When fine-tuned, the relatively small model compete with larger LLMs. In our study, we compare domain-generalist models with those adapted to the medical domain. Since full parameter tuning is costly, many works focus on domain knowledge adaptation by pre-training Shi et al. (2024); Song et al. (2024) or instruction tuning Willard and Louf (2023) with adapters. Training-free approaches encompass chain-of-thought (CoT) Wei et al. (2022); Jeong et al. (2024), self-consistency Wang et al. (2022), and, concerned with lack of grounding resulting in hallucination, recent work introduce RAG methods Li et al. (2024); Wang et al. (2024b); Yu et al. (2023); Munnangi et al. (2024); Wang et al. (2024a); Soong et al. (2023). However, most of these efforts have focused on performance in a particular knowledge paradigm and have lacked a systematic assessment of their performance on structured prediction, which we address in our study.
3 Methodology
Our methodology is designed to answer the following two research questions:
1. How well do Large Language Models (LLMs) perform on structured prediction tasks when provided with unstructured inputs?
2. To what extent can approaches that enhance task knowledge and external knowledge improve their performance?
To answer the first research question, we benchmark LLMs on biomedical text classification and NER datasets, focusing on the “true” zero-shot setting to evaluate models’ parametric knowledge. This reflects real-world scenarios where annotated data is often unavailable due to ad-hoc task requirements, resource limitations and privacy constraints (Giachelle et al., 2021). This leads to what Fries et al. (2022b) describe as “dataset debt”, highlighting issues like inconsistent documentation, lack of domain-specific information except generic entities and difficulties adapting datasets to niche domains. Clinicians face significant time constraints, which limit even few-shot annotations (Xia and Yetisgen-Yildiz, 2012; Wac et al., 2024; Farri et al., 2013). These factors make fine-tuning and few-shot approaches impractical for structured prediction tasks in the biomedical domain, positioning zero-shot methods as a scalable solution for real-world biomedical tasks.
To answer the second question, we compare their zero-shot performance to various methods that aim to enhance task knowledge and external knowledge, while keeping the parametric knowledge static.
@@ 3 Methodology
Table 1 lists our methods. We use Vanilla prompting as the baseline, and enhance it with advanced approaches: chain-of-thought (CoT) Wei et al. (2022) and self-consistency (SC) Wang et al. (2022), along with retrieval-augmented generation (RAG) Lewis et al. (2020) that leverages FAISS with PubMed abstracts and Wikipedia articles, embedding documents via all-MiniLM-L6-v2 Reimers and Gurevych (2019). We also apply constrained decoding Willard and Louf (2023) to enforce structured outputs. For NER, we adopt a two-stage approach: Stage 1 assigns generic entity labels (e.g., “Bodypart”), and Stage 2 refines them to fine-grained labels. Self-consistency is employed in both tasks to aggregate multiple reasoning paths via majority voting.
Technique Details Comments
Vanilla Standard prompting. Baseline for all tasks.
Chain-of-Thought (CoT) Wei et al. (2022) Chain-of-thought reasoning. Effective for QA and logical reasoning. For NER, adapted into a two-stage approach where generic entity names are first induced (e.g., Bodypart), followed by fine-grained labeling.
Self-Consistency (SC) Wang et al. (2022) Majority voting across sampled reasoning paths. Applied in both stages of the two-stage NER approach.
Retrieval-Augmented-Generation (RAG) Lewis et al. (2020) Retrieval-augmented generation using FAISS Douze et al. (2024). Used PubMed Sanyal et al. (2021) and Wikipedia as corpora. PubMed improved performance; Wikipedia degraded performance for medical QA Xiong et al. (2024).
Constrained Decoding Willard and Louf (2023) Restricted outputs to ensure structured extraction. Avoided hallucinations. Ensured span and label consistency in NER tasks.
Table 1: Techniques Summary with Comments and Details. Complete details can be found in Appendix D.
Complete details of our datasets, techniques and methods are described in Appendix D.
4 Evaluation Results
The complete table of results is provided in Table 2. We give an overview of our findings followed by a deeper analysis of the evaluated techniques.
4.1 Overview of results